OpenStudy (anonymous):

The table below represents a linear function f(x) and the equation represents a function g(x): x f(x) x y -1 -9 0 -1 1 7 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). (6 points) Part B: Which function has a greater y-intercept? Justify your answer. (4 points)

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Um, what's function f(x)?

OpenStudy (anonymous):

What You Mean?

OpenStudy (anonymous):

Oh f(x) is where the -9 is an down @iGreen

OpenStudy (anonymous):

x is where the -1 is and down

OpenStudy (anonymous):

g(x) g(x) = 3x - 2

OpenStudy (igreen):

Oh

OpenStudy (igreen):

Okay, I knew about the graph..okay..nvm.

OpenStudy (igreen):

Let's start with f(x). To find the slope, let's take any two points on the line and plug them into the slope formula. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{7+1}{1-0}\) Simplify that

OpenStudy (anonymous):

8 & 0